OpenStudy (anonymous):

The area of a circle with a radius of 15 in., to the nearest tenth is _____. Be sure to include units.

OpenStudy (anonymous):

A = \[A = \pi r^2\] \[A = \pi (15)^2\] \[A= 706\]

OpenStudy (anonymous):

oops. To the nearest tenth is 706.9 in squared

OpenStudy (anonymous):

so just 700 ?

OpenStudy (anonymous):

To the nearest tenth is to one decimal place so it would be 706.9 inches squared

OpenStudy (anonymous):

@HelloKalii what were the answer to this test?